# Information-Theoretic Thresholds for the Alignments of Partially Correlated Graphs

### v1

@@ A.1 Proof of Lemma 1
We first evaluate the moment generating function for paths. Consider a path $P$ of size $\ell$ denoted by $\langle e_{1}e_{2}\dots e_{\ell}\rangle$ as illustrated in Figure 3. For each $i=1,\dots,\ell$, define $A_{i-1}\triangleq{\mathbf{1}_{\left\{{e_{i}\in E(G_{1})}\right\}}}$ and $B_{i}\triangleq{\mathbf{1}_{\left\{{\pi(e_{i})\in E(G_{2})}\right\}}}$. Then $(A_{i},B_{i})\sim\mathrm{Bern}(p,p,\rho)$. By definition (5),
$|P\cap E({\mathcal{H}}_{\pi})|=\sum_{i=1}^{\ell}{\mathbf{1}_{\left\{{e_{i}\in E(G_{1})}\right\}}}{\mathbf{1}_{\left\{{\pi(e_{i})\in E(G_{2})}\right\}}}=\sum_{i=1}^{\ell}A_{i-1}B_{i}.$
@@
For the sake of notational simplicity, we introduce an auxiliary random variable $B_{0}$ that is correlated with $A_{0}$ such that $(A_{0},B_{0})\sim\mathrm{Bern}(p,p,\rho)$. Then
$m_{\ell}\triangleq\mathbb{E}[e^{t|P\cap E({\mathcal{H}}_{\pi})|}]$ $=\mathbb{E}\left[\mathbb{E}\left[\prod_{i=1}^{\ell}e^{tA_{i-1}B_{i}}|B_{0}\dots B_{\ell}\right]\right]=\mathbb{E}\left[\prod_{i=1}^{\ell}\mathbb{E}\left[e^{tA_{i-1}B_{i}}|B_{i-1}B_{i}\right]\right]$
$=\sum_{b_{0},\dots,b_{\ell}\in\{0,1\}}\prod_{i=0}^{\ell}\mathbb{P}[B_{i}=b_{i}]\prod_{i=1}^{\ell}\mathbb{E}\left[e^{tA_{i-1}b_{i}}|B_{i-1}=b_{i-1}\right].$ (16)
Define $M(b_{i-1},b_{i})\triangleq\mathbb{P}[B_{i}=b_{i}]\mathbb{E}\left[e^{tA_{i-1}b_{i}}|B_{i-1}=b_{i-1}\right]$ for $b_{i-1},b_{i}\in\{0,1\}$ and a matrix
@@
where $\bar{p}=1-p$. Recall that $\mathbb{P}[B_{i}=1]=p$. Then we obtain that
$m_{\ell}=\sum_{b_{0},\dots,b_{\ell}\in\{0,1\}}\mathbb{P}[B_{0}=b_{0}]M(b_{0},b_{1})\dots M(b_{\ell-1},b_{\ell})=[\bar{p},p]M^{\ell}\begin{bmatrix}1\\ 1\end{bmatrix}.$
The trace and determinant of $M$ is
@@
Since $D<p_{11}(e^{t}-1)$, the discriminant is $T^{2}-4D>0$. Hence, the matrix $M$ has two distinct eigenvalues denoted by $\lambda_{1}>\lambda_{2}>0$, and the general term of $m_{\ell}$ is
$m_{\ell}=\alpha\lambda_{1}^{\ell}+\beta\lambda_{2}^{\ell}.$ (17)
The coefficients $\alpha$ and $\beta$ can be determined via the first two terms $m_{0}=1$ and $m_{1}$. Then we get
@@
Furthermore, by plugging $m_{1}=1+p^{2}(e^{t}-1)$, we get $T-m_{1}=D$ and thus $m_{1}(T-m_{1})>D$, which is equivalent to $|2m_{1}-T|<\sqrt{T^{2}-4D}$. Therefore, both coefficients $\alpha,\beta\in(0,1)$.
The analysis for cycles follows from similar arguments. Consider a cycle $C$ of size $\ell$ denoted by $[e_{1}\dots e_{\ell}]$ as illustrated in Figure 4. For each $i=1,\dots,\ell$, define $A_{i-1}\triangleq{\mathbf{1}_{\left\{{e_{i}\in E(G_{1})}\right\}}}$ and $B_{i}\triangleq{\mathbf{1}_{\left\{{\pi(e_{i})\in E(G_{2})}\right\}}}$. We also let $B_{0}=B_{\ell}$ for notational simplicity. Then $(A_{i},B_{i})\sim\mathrm{Bern}(p,p,\rho)$ for $i=0,\dots,\ell-1$.
@@
$\tilde{m}_{\ell}\triangleq\mathbb{E}[e^{t|C\cap E({\mathcal{H}}_{\pi})|}]$ $=\sum_{b_{1},\dots,b_{\ell}=b_{0}\in\{0,1\}}\prod_{i=1}^{\ell}\mathbb{P}[B_{i}=b_{i}]\prod_{i=1}^{\ell}\mathbb{E}\left[e^{tA_{i-1}b_{i}}|B_{i-1}=b_{i-1}\right]$
@@
By definition, $\kappa^{{\mathsf{P}}}_{\ell}(t)=\log m_{\ell}$ and $\kappa^{{\mathsf{C}}}_{\ell}(t)=\log\tilde{m}_{\ell}$. To upper bound the cumulants, it suffices to consider $m_{\ell}$ and $\tilde{m}_{\ell}$. In (17), we have $\alpha,\beta\in(0,1)$ and $\lambda_{1}>\lambda_{2}>0$. By monotonicity, it follows that $m_{\ell}\leq\tilde{m}_{\ell}$ and thus
@@
$\log\mathbb{E}\left[e^{t|{\mathcal{E}}\cap E({\mathcal{H}}_{\pi})|}\right]$ $=\sum_{P\in{\mathcal{P}}}\kappa_{|P|}^{{\mathsf{P}}}(t)+\sum_{C\in{\mathcal{C}}}\kappa_{|C|}^{{\mathsf{C}}}(t)$
@@
where the last equality used fact that $|{\mathcal{E}}|=\sum_{P\in{\mathcal{P}}}|P|+\sum_{C\in{\mathcal{C}}}|C|$.
@@
We have two bounds for large $\ell$ in Lemma 1, namely $\kappa_{\ell}^{\mathsf{P}}(t)\leq\kappa_{\ell}^{\mathsf{C}}(t)$ and $\kappa_{\ell}^{\mathsf{C}}(t)\leq\frac{\ell}{2}\kappa_{2}^{\mathsf{C}}(t)$. For the first bound, we apply $\frac{1}{\ell}\log\left(\alpha\lambda_{1}^{\ell}+\beta\lambda_{2}^{\ell}\right)\leq\frac{1}{\ell}\log(\lambda_{1}^{\ell}+\lambda_{2}^{\ell})$, where $10<\beta<\alpha<1,\alpha+\beta=$ 1 and $\lambda_{1}>\lambda_{2}>0$. Consequently, $\lambda_{1}-\frac{\log 2}{\ell}\leq\frac{1}{\ell}\kappa_{\ell}^{\mathsf{P}}(t)\leq\frac{1}{\ell}\kappa_{\ell}^{\mathsf{C}}(t)\leq\lambda_{1}+\frac{\log 2}{\ell}.$ Hence, the first bound is essentially tight for large $\ell$. The second bound, previously used in [WXY22], applies the inequality $\|x\|_{\ell}\leq\|x\|_{2}$, which becomes less tight as $\ell$ increases. Nevertheless, it suffices for our analysis as the probability of long cycles occurring is relatively small.
A.2 Proof of Lemma 2
By Lemma 1,

### v2

@@ B.1 Proof of Lemma 1
We first evaluate the moment generating function for paths. Consider a path $P$ of size $\ell$ denoted by $\langle e_{1}e_{2}\dots e_{\ell}\rangle$ as illustrated in Figure 3. For each $i=1,\dots,\ell$, define $A_{i-1}\triangleq\beta_{e_{i}}(G_{1})$ and $B_{i}\triangleq\beta_{\pi(e_{i})}(G_{2})$. Then $(A_{i},B_{i})\sim\mathrm{Bern}(p,p,\rho)$. By definition (7),
$\beta_{P}(\mathcal{H}_{\pi}^{f})=\sum_{i=1}^{\ell}\beta_{e_{i}}(G_{1})\beta_{\pi(e_{i})}(G_{2})=\sum_{i=1}^{\ell}A_{i-1}B_{i}.$
@@
For notational simplicity, we introduce an auxiliary random variable $B_{0}$ that is correlated with $A_{0}$ such that $(A_{0},B_{0})\sim\mathrm{Bern}(p,p,\rho)$. Then
$m_{\ell}\triangleq\mathbb{E}\left[e^{t\beta_{P}\left(\mathcal{H}^{f}_{\pi}\right)}\right]$ $=\mathbb{E}\left[\mathbb{E}\left[\prod_{i=1}^{\ell}e^{tA_{i-1}B_{i}}|B_{0}\dots B_{\ell}\right]\right]=\mathbb{E}\left[\prod_{i=1}^{\ell}\mathbb{E}\left[e^{tA_{i-1}B_{i}}|B_{i-1},B_{i}\right]\right]$
$=\sum_{b_{0},\dots,b_{\ell}\in\{0,1\}}\prod_{i=0}^{\ell}\mathbb{P}\left[B_{i}=b_{i}\right]\prod_{i=1}^{\ell}\mathbb{E}\left[e^{tA_{i-1}b_{i}}|B_{i-1}=b_{i-1}\right].$ (38)
Define $M(b_{i-1},b_{i})\triangleq\mathbb{P}\left[B_{i}=b_{i}\right]\mathbb{E}\left[e^{tA_{i-1}b_{i}}|B_{i-1}=b_{i-1}\right]$ for $b_{i-1},b_{i}\in\{0,1\}$ and a matrix
@@
where $\bar{p}\triangleq 1-p$. Recall that $\mathbb{P}\left[B_{i}=1\right]=p$. Then, we obtain that
$m_{\ell}=\sum_{b_{0},\dots,b_{\ell}\in\{0,1\}}\mathbb{P}\left[B_{0}=b_{0}\right]M(b_{0},b_{1})\dots M(b_{\ell-1},b_{\ell})=[\bar{p},p]M^{\ell}\begin{bmatrix}1\\ 1\end{bmatrix}.$
The trace and determinant of $M$ are given by
@@
Since $D<p_{11}(e^{t}-1)$, the discriminant is $T^{2}-4D>0$. Hence, the matrix $M$ has two distinct eigenvalues denoted by $\lambda_{1}>\lambda_{2}$. Since $\lambda_{1}+\lambda_{2}=T>0$ and $\lambda_{1}\lambda_{2}=D>0$, we have $\lambda_{1}>\lambda_{2}>0$, and the general term of $m_{\ell}$ is
$m_{\ell}=\alpha_{1}\lambda_{1}^{\ell}+\alpha_{2}\lambda_{2}^{\ell}.$ (39)
The coefficients $\alpha_{1}$ and $\alpha_{2}$ can be determined via the first two terms $m_{0}=1$ and $m_{1}$. Then we get
@@
Furthermore, by plugging $m_{1}=1+p^{2}(e^{t}-1)$, we get $T-m_{1}=D$ and thus $m_{1}(T-m_{1})>D$, which is equivalent to $|2m_{1}-T|<\sqrt{T^{2}-4D}$. Therefore, both coefficients $\alpha_{1},\alpha_{2}\in(0,1)$.
The analysis for cycles follows from similar arguments. Consider a cycle $C$ of size $\ell$ denoted by $[e_{1}\dots e_{\ell}]$ as illustrated in Figure 4. For each $i=1,\dots,\ell$, define $A_{i-1}\triangleq\beta_{e_{i}}(G_{1})$ and $B_{i}\triangleq\beta_{\pi(e_{i})}(G_{2})$. We also let $B_{0}=B_{\ell}$ for notational simplicity. Then $(A_{i},B_{i})\sim\mathrm{Bern}(p,p,\rho)$ for $i=0,\dots,\ell-1$.
@@
$\tilde{m}_{\ell}\triangleq\mathbb{E}\left[e^{t\beta_{C}\left(\mathcal{H}^{f}_{\pi}\right)}\right]$ $=\sum_{b_{1},\dots,b_{\ell}=b_{0}\in\{0,1\}}\prod_{i=1}^{\ell}\mathbb{P}\left[B_{i}=b_{i}\right]\prod_{i=1}^{\ell}\mathbb{E}\left[e^{tA_{i-1}b_{i}}|B_{i-1}=b_{i-1}\right]$
@@
By definition, $\kappa^{{\mathsf{P}}}_{\ell}(t)=\log m_{\ell}$ and $\kappa^{{\mathsf{C}}}_{\ell}(t)=\log\tilde{m}_{\ell}$. To upper bound the cumulants, it suffices to consider $m_{\ell}$ and $\tilde{m}_{\ell}$. In (39), we have $\alpha_{1},\alpha_{2}\in(0,1)$ and $\lambda_{1}>\lambda_{2}>0$. By monotonicity, it follows that $m_{\ell}\leq\tilde{m}_{\ell}$ and thus
@@
$\log\mathbb{E}\left[e^{t\beta_{\mathcal{E}}(\mathcal{H}_{\pi}^{f})}\right]$ $=\sum_{P\in{\mathcal{P}}}\kappa_{|P|}^{{\mathsf{P}}}(t)+\sum_{C\in{\mathcal{C}}}\kappa_{|C|}^{{\mathsf{C}}}(t)$
@@
where the last equality uses fact that $|{\mathcal{E}}|=\sum_{P\in{\mathcal{P}}}|P|+\sum_{C\in{\mathcal{C}}}|C|$.
@@
We have two bounds for large $\ell$ in the Erdős-Rényi model, namely $\kappa_{\ell}^{\mathsf{P}}(t)\leq\kappa_{\ell}^{\mathsf{C}}(t)$ and $\kappa_{\ell}^{\mathsf{C}}(t)\leq\frac{\ell}{2}\kappa_{2}^{\mathsf{C}}(t)$. For the first bound, we apply $\frac{1}{\ell}\log\left(\alpha_{1}\lambda_{1}^{\ell}+\alpha_{2}\lambda_{2}^{\ell}\right)\leq\frac{1}{\ell}\log(\lambda_{1}^{\ell}+\lambda_{2}^{\ell})$, where $10<\alpha_{2}<\alpha_{1}<1,\alpha_{1}+\alpha_{2}=$ 1 and $\lambda_{1}>\lambda_{2}>0$. Consequently, $\lambda_{1}-\frac{\log 2}{\ell}\leq\frac{1}{\ell}\kappa_{\ell}^{\mathsf{P}}(t)\leq\frac{1}{\ell}\kappa_{\ell}^{\mathsf{C}}(t)\leq\lambda_{1}+\frac{\log 2}{\ell}.$ Hence, the first bound is essentially tight for large $\ell$. The second bound, previously used in [WXY22], applies the inequality $\|x\|_{\ell}\leq\|x\|_{2}$, which becomes less tight as $\ell$ increases. Nevertheless, it suffices for our analysis as the probability of long cycles occurring is relatively small.
Gaussian Wigner Model, Part 1
In this part, we focus on the Gaussian Wigner model with $f(x,y)=xy$. The lower-order cumulants can be calculated directly:
$\kappa_{1}^{\mathsf{C}}(t)$ $=-\frac{1}{2}\log(1-2t\rho-t^{2}(1-\rho^{2})),\quad\kappa_{1}^{\mathsf{P}}(t)=-\frac{1}{2}\log\left(1-t^{2}\right),$ (41)
$\kappa_{2}^{\mathsf{C}}(t)$ $=-\frac{1}{2}\log\left(1-2t^{2}(1+\rho^{2})+t^{4}(1-\rho^{2})^{2}\right).$ (42)
We first evaluate the moment generating function for paths. Consider a path $P$ of size $\ell$ denoted by $\langle e_{1}e_{2}\dots e_{\ell}\rangle$ as illustrated in Figure 3. For each $i=1,\dots,\ell$, define $A_{i-1}\triangleq\beta_{e_{i}}(G_{1})$ and $B_{i}\triangleq\beta_{\pi(e_{i})}(G_{2})$. Then $(A_{i},B_{i})\sim\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}\right)$. By definition (7),
$\beta_{P}(\mathcal{H}^{f}_{\pi})=\sum_{i=1}^{\ell}\beta_{e_{i}}(G_{1})\beta_{\pi(e_{i})}(G_{2})=\sum_{i=1}^{\ell}A_{i-1}B_{i}.$
For the sake of notational simplicity, we introduce an auxiliary variable $B_{0}$ that is correlated with $A_{0}$ such that $(A_{0},B_{0})\sim\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}\right)$. Then
$m_{\ell}\triangleq\mathbb{E}\left[e^{t\beta_{P}\left(\mathcal{H}^{f}_{\pi}\right)}\right]$ $=\mathbb{E}\left[\mathbb{E}\left[\prod_{i=0}^{\ell-1}e^{tA_{i}B_{i+1}}|B_{0}\dots B_{\ell}\right]\right]=\mathbb{E}\left[\prod_{i=0}^{\ell-1}\mathbb{E}\left[e^{tA_{i}B_{i+1}}|B_{i},B_{i+1}\right]\right]$
$=\mathbb{E}\left[\prod_{i=0}^{\ell-1}\exp\left(t\rho B_{i}B_{i+1}+\frac{1}{2}t^{2}(1-\rho^{2})B_{i+1}^{2}\right)\right],$ (43)
where the last equality follows from $A_{i}|B_{i}\sim\mathcal{N}(\rho B_{i},1-\rho^{2})$ and $\mathbb{E}\left[\exp(tZ)\right]=\exp\left(t\mu+t^{2}\nu^{2}/2\right)$ for $Z\sim\mathcal{N}(\mu,\nu^{2})$.
Define $\mathbf{B}_{\ell}\triangleq[B_{0},B_{1},\cdots,B_{\ell}]$ and a matrix
$\mathbf{W}_{\ell}\triangleq\begin{bmatrix}1&-t\rho&0&\cdots&0\\ -t\rho&1-t^{2}(1-\rho^{2})&-t\rho&\cdots&0\\ 0&-t\rho&1-t^{2}(1-\rho^{2})&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&-t\rho\\ 0&0&\cdots 0&-t\rho&1-t^{2}(1-\rho^{2})\end{bmatrix},$
where $(\mathbf{W}_{\ell})_{00}=1,(\mathbf{W}_{\ell})_{ii}=1-t^{2}(1-\rho^{2})$ for any $1\leq i\leq\ell$ and $(\mathbf{W}_{\ell})_{ij}=-t\rho$ for any $|i-j|=1$ with $0\leq i,j\leq\ell$. By (43), we have
$m_{\ell}$ $=\mathbb{E}\left[\prod_{i=0}^{\ell-1}\exp\left(t\rho B_{i}B_{i+1}+\frac{1}{2}t^{2}(1-\rho^{2})B_{i+1}^{2}\right)\right]$
$=\int\cdots\int\left(\frac{1}{\sqrt{2\pi}}\right)^{\ell+1}\exp\left(-\frac{1}{2}\sum_{i=0}^{\ell}B_{i}^{2}\right)\exp\left(\sum_{i=0}^{\ell-1}\left(t\rho B_{i}B_{i+1}+\frac{1}{2}t^{2}(1-\rho^{2})B_{i+1}^{2}\right)\right)\,\mathrm{d}B_{0}\cdots\mathrm{d}B_{\ell}$
$=\int\cdots\int\left(\frac{1}{\sqrt{2\pi}}\right)^{\ell+1}\exp\left(-\frac{1}{2}\mathbf{B}_{\ell}^{\top}\mathbf{W}_{\ell}\mathbf{B}_{\ell}\right)\,\mathrm{d}B_{0}\cdots\mathrm{d}B_{\ell}.$ (44)
If $\mathbf{W}_{\ell}$ is positive definite, then $m_{\ell}=\det\left(\mathbf{W}_{\ell}\right)^{-1/2}$ by (44). We then prove $\mathbf{W}_{\ell}$ is positive definite and compute the explicit formula of $\det\left(\mathbf{W}_{\ell}\right)$. Expanding the last column of $\mathbf{W}_{\ell}$ yields that
$\det(\mathbf{W}_{\ell})=\left(1-t^{2}(1-\rho^{2})\right)\det(\mathbf{W}_{\ell-1})-t^{2}\rho^{2}\det(\mathbf{W}_{\ell-2}),\text{ for any }\ell\geq 2.$
Therefore, the general term of $\det(\mathbf{W}_{\ell})$ is determined by
$\det(\mathbf{W}_{\ell})=\alpha_{1}\lambda_{1}^{\ell}+\alpha_{2}\lambda_{2}^{\ell},$
where $\lambda_{1}>\lambda_{2}$ are two roots of equation $x^{2}-\left(1-t^{2}(1-\rho^{2})\right)x+t^{2}\rho^{2}=0$. The coefficients $\alpha_{1}$ and $\alpha_{2}$ can be determined via the first two terms $\det(\mathbf{W}_{0})=1$ and $\det(\mathbf{W}_{1})=1-t^{2}$. Specifically,
$\alpha_{1}={\frac{1}{2}+\frac{1-t^{2}(1+\rho^{2})}{2\sqrt{(1-t^{2}(1-\rho^{2}))^{2}-4t^{2}\rho^{2}}}},\quad\alpha_{2}={\frac{1}{2}-\frac{1-t^{2}(1+\rho^{2})}{2\sqrt{(1-t^{2}(1-\rho^{2}))^{2}-4t^{2}\rho^{2}}}}.$
Since $01-t^{2}(1-\rho^{2})-2t\rho=(1-t\rho-t)(1-t\rho+t)>$ 0 by $t<\frac{1}{1+\rho}$, we obtain $(1-t^{2}(1-\rho^{2}))^{2}-4t^{2}\rho^{2}>0$, and hence
$\frac{1-t^{2}(1+\rho^{2})}{\sqrt{(1-t^{2}(1-\rho^{2}))^{2}-4t^{2}\rho^{2}}}=\sqrt{\frac{(1-t^{2}(1-\rho^{2}))^{2}-4t^{2}\rho^{2}+4t^{4}\rho^{2}}{(1-t^{2}(1-\rho^{2}))^{2}-4t^{2}\rho^{2}}}>1,$
which implies that $\alpha_{1}>1$ and $\alpha_{2}<0$. Since $\lambda_{1}+\lambda_{2}=1-t^{2}(1-\rho^{2})>0$ by $t<\frac{1}{1+\rho}$ and $\lambda_{1}\lambda_{2}=t^{2}\rho^{2}>0$ and the discriminant $(1-t^{2}(1-\rho^{2}))^{2}-4t^{2}\rho^{2}>0$ , we have $\lambda_{1}>\lambda_{2}>0$. Therefore,
$\det(\mathbf{W}_{\ell})=\alpha_{1}\lambda_{1}^{\ell}+\alpha_{2}\lambda_{2}^{\ell}=\alpha_{1}(\lambda_{1}^{\ell}-\lambda_{2}^{\ell})+\lambda_{2}^{\ell}>0,\text{ for any }\ell\geq 0,$
which implies that $\mathbf{W}_{\ell}$ is positive definite by Sylvester’s criterion (see, e.g., [HJ12, Theorem 7.2.5]). Denote $\mathbf{W}_{\ell}=\mathbf{J}_{\ell}^{\top}\mathbf{J}_{\ell}$ with $\mathbf{J}_{\ell}\in\mathbb{R}^{(\ell+1)\times(\ell+1)}$. Then, by (44),
$m_{\ell}$ $=\int\cdots\int\left(\frac{1}{\sqrt{2\pi}}\right)^{\ell+1}\exp\left(-\frac{1}{2}\mathbf{B}_{\ell}^{\top}\mathbf{W}_{\ell}\mathbf{B}_{\ell}\right)\,\mathrm{d}B_{0}\cdots\mathrm{d}B_{\ell}$
$=\int\cdots\int\left(\frac{1}{\sqrt{2\pi}}\right)^{\ell+1}\exp\left(-\frac{1}{2}\mathbf{B}_{\ell}^{\top}\mathbf{J}_{\ell}^{\top}\mathbf{J}_{\ell}\mathbf{B}_{\ell}\right)\,\mathrm{d}B_{0}\cdots\mathrm{d}B_{\ell}=\left[\det\left(\mathbf{J}_{\ell}\right)\right]^{-1}=\left[\det\left(\mathbf{W}_{\ell}\right)\right]^{-1/2}.$
The analysis for cycles follows from similar arguments. Consider a cycle $C$ of size $\ell$ denoted by $[e_{1}\dots e_{\ell}]$ as illustrated in Figure 4. For each $i=1,\dots,\ell$, define $A_{i-1}\triangleq\beta_{e_{i}}(G_{1})$ and $B_{i}\triangleq\beta_{\pi(e_{i})}(G_{2})$. We also let $B_{0}=B_{\ell}$ for notational simplicity. Then $(A_{i},B_{i})\sim\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}\right)$ for $i=0,\dots,\ell-1$. Recall that $\lambda_{1}+\lambda_{2}=1-t^{2}(1-\rho^{2})$ and $\lambda_{1}\lambda_{2}=t^{2}\rho^{2}$. Following a similar argument as (43), we have